OpenStudy (anonymous):

Which number serves as a counterexample to the statement below? All positive integers are divisible by 2 or 3.......PLEASE EXPLAIN :]

OpenStudy (regan):

11

myininaya (myininaya):

lol you beat me i was going to say 13

OpenStudy (regan):

any prime that isnt 2 or 3

myininaya (myininaya):

there are also some composite numbers that dont have factors 2 and 3 like 77

OpenStudy (anonymous):

the answer choices are 100, 57,30, or 25...please explain how you get the answer

OpenStudy (regan):

hmmmm true very interesting

myininaya (myininaya):

100/2=50 57/3=19 30/3=10 25 is not divisible by either

myininaya (myininaya):

25/2 is not an integer so 25 is not divisible by 2 25/3 is not an integer so 25 is not divisible by 3

myininaya (myininaya):

cool?

myininaya (myininaya):

here regan i will give you some medals

OpenStudy (anonymous):

Thankk Yhuu Guys ! <3

myininaya (myininaya):

saffire i will give you a medal if you tell me if 27 is divisible by 3

OpenStudy (regan):

thank you : ) if you describe the whole set ill give you one

OpenStudy (anonymous):

yes it is divisible by 3

myininaya (myininaya):

vg saffire u get a medal lol

OpenStudy (anonymous):

HaaHaa Thank Yhuu

myininaya (myininaya):

{x is an element of the Z|2divides x and 3 divides x} the set has been described lol

myininaya (myininaya):

or*